OpenStudy (thefluffmuncher):

simplify squareroot of 1008

OpenStudy (anonymous):

1008 = 2 * 2 * 2 * 2 * 3 * 3 * 7 so its (2 ^ 4) * (3 ^ 2) * 7 which is equal to 12 * sort(7)

OpenStudy (anonymous):

12 * sqrt(7)

OpenStudy (anonymous):

\[\sqrt{1008} = 12 * \sqrt{7}\]